Jenny invested her savings in two investment funds.The amount she invested in fund A was 3 times as much as the amount she invested in fund B. Fund A returned a 3% profit and fund B returned a 2% profit. How much did she invest in fund B, if the total profit from the two funds together was $990?
Answer by Theo(13342)

let a = the amount invested in fund A.
let b = the amount invested in fund B.
the amount invested in fund A is 3 * the amount invested in fund B.
a = 3b
fund A returns 3% profit.
fund B returns 2% profit.
total profit is 990
.03a + .02b = 990
since a = 3b, replace a with 3b to get:
.03 * (3b) + .02b = 990
simplify to get .09b + .02b = 990
combine like terms to get .11b = 990
solve for b to get b = 990 / .11 = 9000
since a = 3b, then a = 27000
.03 * a + .02 *b = 810 + 180 = 990
solution looks good.
solution is that the amount invested in fund B is 9000.